# Tensor polarization and the dissipative damping of vector meson in QCD Medium

### v5

@@ Abstract
To obtain useful formulations for these quantities, the theory is developed in the following context: vector mesons with momentum $\tilde{p}^{\mu}=(\varepsilon_{\bm{p}},\bm{p})$ at density $n(\varepsilon_{u})$ propagate in a locally equilibrated hydrodynamic medium, where their tensor polarization ${\cal T}^{\mu\nu}$ can be characterized by hydrodynamic variables such as inverse temperature $\beta=1/T$, the flow velocity $u^{\mu}(x)$, symmetric flow gradients $\xi_{\lambda\gamma}\equiv\beta^{-1}\partial_{(\lambda}(\beta u)_{\gamma)}$, etc. To study general polarization phenomena in this context, we have developed a comprehensive theoretical framework “Zubarev response approach”, which is detailed in “supplementary materials” and in our long theory paper Li and Liu (2025). Especially, in the long paper, we illustrate how our “Zubarev response approach”, rooted rigorously in thermal field theory, can systematically calculate various polarization phenomena, including the dissipative and non-dissipative contribution on equal footing, and lay out the path to incorporating non-perturbative methods, such as lattice methods and the functional renormalization group.
@@
where the spectral function $A(\omega,\bm{p})$ controls the dissipative damping of a particle’s propagation as $D(t,x)\propto\int d\omega\,e^{-i\omega t+\bm{p}\cdot x}A(\omega,\bm{p})$. The projectors in Eq. (1) are defined as $\tilde{\Delta}^{\mu\nu}=-\eta^{\mu\nu}+\tilde{p}^{\mu}\tilde{p}^{\nu}/m^{2}\;$ and $\tilde{\Delta}^{\langle\mu}_{\;\lambda}\tilde{\Delta}^{\nu\rangle}_{\;\gamma}=\tilde{\Delta}^{(\mu}_{\;\lambda}\tilde{\Delta}^{\nu)}_{\;\gamma}-\tilde{\Delta}^{\mu\nu}\tilde{\Delta}_{\lambda\gamma}/3$. The second line of Eq.(Tensor polarization and the dissipative damping of vector meson in QCD Medium) is employed a quasi-particle approximation, where the $\Delta\varepsilon_{\bm{p}}$ is the in-medium energy shift and is T-even. The $\Gamma_{\bm{p}}$ is the width of the particle, i.e. twice damping rate, of the particles in-medium, which is the T-odd and is the source of dissipation in this calculation. As shown, the $\alpha_{\rm sh}$ is an odd power of $\Gamma_{\bm{p}}$ , so it is T-odd and dissipative.
@@
By correlating beam energies and centralities with spectral properties, such as mass shift or width, we can generate spin alignments with rich behavior. In Fig. 1, the $2^{\text{nd}}$ figure demonstrates that a change of less than 40 MeV (ansatzes in figures) in the in-medium $\phi$ mass with beam energy mimics the observed beam energy dependence of $\rho_{00}$ for $\phi$ and $K^{*}$. The out-plane $\rho_{00}$ is automatically separated from the in-plane one. Although the splitting size depends on the details of the hydrodynamic simulation, the general hierarchy is robust and resembles that observed in experiments. We employ different illustrative ansatzes for the spectral properties of $K^{*}$ and $\phi$, reflecting their different in-medium interactions and freezeout stages: $K^{*}$ freezes out in the late hadronic phase (positive mass-shifts Ilner et al. (2017)), while $\phi$ is assumed to freeze out earlier (QGP or early hadronic phase).The corresponding $p_{T}$ dependence of $\rho_{00}$ is shown in the $3^{\text{rd}}$ figure, which can accommodate a $p_{T}$-dependent sign-flipping behavior. Correlating spectral properties with centrality ($4^{\text{th}}$ figure) can generate a sign-flipping $\rho_{00}$ dependence, mimicking those observed in experiments.
@@
Quark-meson model coefficients.— To further justify physically motivated coefficients and obtain the model-independent insights, we have done a microscopic calculation on the coefficients and used it to calculate the spin alignment in this section. We first evaluate the coefficients $\alpha_{\text{sh}}$ numerically using Eq. (Tensor polarization and the dissipative damping of vector meson in QCD Medium) for $\phi$ meson with in-medium spectral functions obtained via a 1-loop quark meson model as discussed in “supplementary materials” (similar ideas in Ref. Dong et al. (2023)). With coupling $g_{\phi}=2$, these coefficients are plotted in Fig. 2 for the cases with the effective strange quark masses $m_{s}=\{0.3,0.42,0.5\}$ GeV, which are relevant mass scales of strangeness at the QGP phase, mix phase, and hadronic phase. In the hadronic phase, this mass represents a effective in-medium kaon mass (see discussion later).
@@
The $2^{\rm nd}$ and $3^{\rm rd}$ panels in Fig. 2 illustrate the $p_{T}$ dependence and the “centrality dependence” of the spin alignment from the QM model. For these figures, we actually include all new effects (splitting, zeroth order) discussed in “supplementary materials” and we confirm that the SITP effect is the dominant one in this setup. As shown, microscopic and more complete calculation can recover the sign-flipping $p_{T}$ and centrality dependence illustrated Fig. 1. While the $p_{T}$ dependence has already been discussed, the correlation explaining the “centrality dependence” is illustrated in the $4^{\rm th}$ panel that we utilize hydrodynamic simulations with a different (spin) freezeout temperature $T_{\rm fz}$ correlated with centrality, which is then corrected with $m_{s}(T_{\rm fz})$ and those $\alpha_{\rm sh}$ coefficients.
@@
The prescription that (spin) freezeout $T$ increases with centrality (more peripheral) resembles the trend in kinetic freeze-out Adamczyk et al. (2017b). At larger centrality, the lifetime of the hadronic phase is much shorter Knospe et al. (2016, 2021); Adam et al. (2017); Acharya et al. (2021) (around 7 fm at 0% to 2 fm at 60% at top RHIC Knospe et al. (2016)). The shorter hadronic phase, which inhibits re-equilibration in a later, lower-temperature stage, effectively raises the “freezeout temperature”. Meanwhile, the mass-dropping scheme employed for the effective s-quark (or other constituents, such as kaons) is mostly motivated by chiral symmetry restoration physics Buballa (2005); Li (2016) and emerging low-lying QGP collective modes Liu and Rapp (2018) that might be hundreds of MeV below the condensate mass. On the other hand, for those who have concerns about the idea of the strange-masses-melting scenario, we can go back to the Fig. 1, where the mass-shifts and width changes required is less than 100 MeV. This is not a strong requirement or unreasonable assumption. Collisional interactions in the QGP phase can generate negative mass shifts as large as a few hundred MeV Liu and Rapp (2018), while radiative ($\phi\rightarrow\bar{K}K\,\text{or}\,\bar{s}s$ type) interactions generate positive mass-shifts as shown in our QM model or Haglin and Gale (1994). Therefore, if the interaction type transitions from collisional to radiative type as it crosses the phase boundary, we can expect a sign-flipping in mass shift, leading to sign-flipping in spin alignment.
@@
Besides published data, preliminary experimental data Xi (2024) suggest a strong rapidity dependence at lower energies. While the SITP effect is largely boost-invariant at mid-rapidity, this dependence can be accounted for by including the zeroth-order term with coefficient $\alpha_{0}$ (see Eq. (I.1) in the Supplemental Material). Tuning both $\alpha_{0}$ and $\alpha_{\rm sh}$ can accommodate this new feature without qualitatively altering the features shown in Fig. 1. A large rapidity dependence would imply a $\phi$ meson with a longitudinal/transverse mass splitting at the order of 100 MeV (see large splitting in other particles Yeo et al. (2024)), a possible case for a strongly-coupled medium (see Fig. 5 in Supplemental Material).
@@
Meanwhile, we also demonstrate that a few-tens-MeV change in mass-shifts and width can significantly change the transport coefficients associated with the above gradients, generating distinguishable behaviors in experimental observables and making spin alignment serve as a promising “spectrometer” candidate for vector mesons
In the future, we will further improve the theoretical derivation with more sophisticated microscopic calculations of the in-medium spectral properties of vector mesons, and we will also explore additional contributions beyond the leading orders Li and Liu (2025). Meanwhile, we need to perform more systematic phenomenological studies to quantitative compare the experimental data, and to accommodate non-equilibrium effects and other novel contributions proposed in Refs. Sheng et al. (2023); Müller and Yang (2021). Additionally, the presence of the newly discovered SITP effect is robust against variations in the details of the interactions and is universal in both relativistic and non-relativistic cases. Therefore, it will be interesting to investigate this effect experimentally in low-energy physics, such as plasma physics or cold atom physics Gregory et al. (2017).
@@ Acknowledgements.
Acknowledgments We gratefully acknowledge the valuable contribution from Yi Yin, who definitely deserves a key position in the author list, but waive the authorship generously. We extend special thanks to Xiangyu Wu for sharing the most recent CLVisc code and providing helpful guidance for our simulations at various beam energies in later versions. We are also especially grateful to Longgang Pang for providing hydrodynamic profiles for our early numerical studies. We also thank Youyu Li, Xinxiang Liu, and Tan Luo for engaging in valuable discussions and providing guidance on hydrodynamic simulations. We thank Hengtong Ding, Xiaojian Du, Aiqiang Guo, Koichi Hattori, Min He, Che-Ming Ko, Yutie Liang, Shu Lin, Rapp Ralf, Kaijia Sun, Subhash Singha, Shuzhe Shi, Yifeng Sun, Aihong Tang, Biaogang Wu, Dilun Yang, Wenbin Zhao for valuable discussions. SL is supported by NSFC No. 12205090. Fundamental Research Funds for the central Universities. FL is supported by NSFC No. 12105129.

### v6

@@ Abstract
To obtain useful formulations for these quantities, the theory is developed in the following context: vector mesons with momentum $\tilde{p}^{\mu}=(\varepsilon_{\bm{p}},\bm{p})$ at density $n(\varepsilon_{u})$ propagate in a locally equilibrated hydrodynamic medium, where their tensor polarization ${\cal T}^{\mu\nu}$ can be characterized by hydrodynamic variables such as inverse temperature $\beta=1/T$, the flow velocity $u^{\mu}(x)$, symmetric flow gradients $\xi_{\lambda\gamma}\equiv\beta^{-1}\partial_{(\lambda}(\beta u)_{\gamma)}$, etc. To study general polarization phenomena in this context, we have developed a comprehensive theoretical framework “Zubarev response approach”, which is detailed in Supplementary Materials (Supp) and in our long theory paper Li and Liu (2025). Especially, in the long paper, we illustrate how our “Zubarev response approach”, rooted rigorously in thermal field theory, can systematically calculate various polarization phenomena, including the dissipative and non-dissipative contribution on equal footing, and lay out the path to incorporating non-perturbative methods, such as lattice methods and the functional renormalization group.
@@
where the spectral function $A(\omega,\bm{p})$ controls the dissipative damping of a particle’s propagation as $D(t,x)\propto\int d\omega\,e^{-i\omega t+\bm{p}\cdot x}A(\omega,\bm{p})$. The projectors in Eq. (1) are defined as $\tilde{\Delta}^{\mu\nu}=-\eta^{\mu\nu}+\tilde{p}^{\mu}\tilde{p}^{\nu}/m^{2}\;$ and $\tilde{\Delta}^{\langle\mu}_{\;\lambda}\tilde{\Delta}^{\nu\rangle}_{\;\gamma}=\tilde{\Delta}^{(\mu}_{\;\lambda}\tilde{\Delta}^{\nu)}_{\;\gamma}-\tilde{\Delta}^{\mu\nu}\tilde{\Delta}_{\lambda\gamma}/3$. The second line of Eq.(Tensor polarization and the dissipative damping of vector meson in QCD Medium) employs a quasi-particle approximation, where the $\Delta\varepsilon_{\bm{p}}$ is the in-medium energy shift and is T-even. The $\Gamma_{\bm{p}}$ is the width of the particle, i.e. twice damping rate, of the particles in-medium, which is the T-odd and is the source of dissipation in this calculation. As shown, the $\alpha_{\rm sh}$ is an odd power of $\Gamma_{\bm{p}}$ , so it is T-odd and dissipative.
@@
By correlating beam energies and centralities with spectral properties, such as mass shift or width, we can generate spin alignments with rich behavior. In Fig. 1, the $2^{\text{nd}}$ figure demonstrates that a change of less than 40 MeV (ansatzes in figures) in the in-medium $\phi$ mass with beam energy mimics the observed beam energy dependence of $\rho_{00}$ for $\phi$ and $K^{*}$. The out-plane $\rho_{00}$ is automatically separated from the in-plane one. Although the splitting size depends on the details of the hydrodynamic simulation, the general hierarchy is robust and resembles that observed in experiments. We employ different illustrative ansatzes for the spectral properties of $K^{*}$ and $\phi$, reflecting their different in-medium interactions and freezeout stages: $K^{*}$ freezes out in the late hadronic phase (positive mass-shifts Ilner et al. (2017)), while $\phi$ is assumed to freeze out earlier (QGP or early hadronic phase).The corresponding $p_{T}$ dependence of $\rho_{00}$ is shown in the $3^{\text{rd}}$ figure, which can accommodate a $p_{T}$-dependent sign-flipping behavior. Correlating spectral properties with $\sqrt[3]{dN/d\eta}$, fitted linearly with centrality Adare et al. (2016) ($4^{\text{th}}$ figure, more in Fig. 6 Supp), can generate a sign-flipping $\rho_{00}$ dependence, mimicking those observed in experiments.
@@
Quark-meson model coefficients.— To further justify physically motivated coefficients and obtain the model-independent insights, we have done a microscopic calculation on the coefficients and used it to calculate the spin alignment in this section. We first evaluate the coefficients $\alpha_{\text{sh}}$ numerically using Eq. (Tensor polarization and the dissipative damping of vector meson in QCD Medium) for $\phi$ meson with in-medium spectral functions obtained via a 1-loop quark meson model as discussed in Supp (similar ideas in Ref. Dong et al. (2023)). With coupling $g_{\phi}=2$, these coefficients are plotted in Fig. 2 for the cases with the effective strange quark masses $m_{s}=\{0.3,0.42,0.5\}$ GeV, which are relevant mass scales of strangeness at the QGP phase, mix phase, and hadronic phase. In the hadronic phase, this mass represents a effective in-medium kaon mass (see discussion later).
@@
The $2^{\rm nd}$ and $3^{\rm rd}$ panels in Fig. 2 illustrate the $p_{T}$ dependence and the centrality dependence of the spin alignment from the QM model. For these figures, we actually include all new effects (splitting, zeroth order) discussed in Supp and we confirm that the SITP effect is the dominant one in this setup. As shown, microscopic and more complete calculation can recover the sign-flipping $p_{T}$ and centrality dependence illustrated Fig. 1. While the $p_{T}$ dependence has already been discussed, the“centrality-freezeout correlation” is illustrated in the $4^{\rm th}$ panel that we utilize hydrodynamic simulations with a different (spin) freezeout temperature $T_{\rm fz}$ correlated with centrality, which is then corrected with $m_{s}(T_{\rm fz})$ and those $\alpha_{\rm sh}$ coefficients.
@@
The plotted “centrality-freezeout correlation”, where $T_{fz}$ increases with centrality (more peripheral), resembles the trend in kinetic freeze-out Adamczyk et al. (2017b). The hadronic phase lifetime is largely linear with $\sqrt[3]{dN/d\eta}$ Knospe et al. (2016, 2021); Adam et al. (2017); Acharya et al. (2021). Since $\sqrt[3]{dN/d\eta}$ also decreases linearly with centrality (see Fig.  6), the lifetime drops from 7 fm (0%) to 2 fm (60%) at top RHIC. At large centrality, the shorter hadronic phase, which inhibits re-equilibration in a later, lower-temperature stage, effectively raises the “freezeout temperature”. Meanwhile, the mass-dropping scheme employed for the effective s-quark (or other constituents, such as kaons) is mostly motivated by chiral symmetry restoration physics Buballa (2005); Li (2016) and emerging low-lying QGP collective modes Liu and Rapp (2018) that might be hundreds of MeV below the condensate mass. On the other hand, for those who have concerns about the idea of the strange-masses-melting scenario, we can go back to the Fig. 1, where the mass-shifts and width changes required is less than 100 MeV. This is not a strong requirement or unreasonable assumption. Collisional interactions in the QGP phase can generate negative mass shifts as large as a few hundred MeV Liu and Rapp (2018), while radiative ($\phi\rightarrow\bar{K}K\,\text{or}\,\bar{s}s$ type) interactions generate positive mass-shifts as shown in our QM model or Haglin and Gale (1994). Therefore, if the interaction type transitions from collisional to radiative type as it crosses the phase boundary, we can expect a sign-flipping in mass shift, leading to sign-flipping in spin alignment.
@@
Besides published data, preliminary experimental data Wilks (2025) suggest a strong rapidity dependence at lower energies. While the SITP effect is largely boost-invariant at mid-rapidity, this dependence can be accounted for by including the zeroth-order term with coefficient $\alpha_{0}$ (see Eq. (I.1) Supp). Tuning both $\alpha_{0}$ and $\alpha_{\rm sh}$ can accommodate this new feature without qualitatively altering the features shown in Fig. 1. A large rapidity dependence would imply a $\phi$ meson with a longitudinal/transverse mass splitting at the order of 100 MeV (see large splitting in other particles Yeo et al. (2024)), a possible case for a strongly-coupled medium (see Fig. 5 Supp).
@@
Meanwhile, we also demonstrate that a few-tens-MeV change in mass-shifts and width can significantly change the transport coefficients associated with the above gradients, generating distinguishable behaviors in experimental observables and making spin alignment serve as a promising “spectrometer” candidate for vector mesons.
To fully realize this “spectrometer” blueprint, we need to investigate additional contributions beyond the leading orders to refine the relation between the spectral function and spin alignment. This effort also requires more sophisticated studies of in-medium spectral properties, alongside further improved phenomenology and simulations. Investigating other non-equilibrium effects and novel contributions proposed in Refs. Sheng et al. (2023); Müller and Yang (2021) is also necessary to understand their quantitative impact. Finally, as a separate point, the newly discovered SITP effect is robust against variations in interaction details and is universal in both relativistic and non-relativistic cases, making it interesting to investigate experimentally in low-energy physics, such as plasma or cold atom physics Gregory et al. (2017).
@@ Acknowledgements.
Acknowledgments We gratefully acknowledge the valuable contribution from Yi Yin, who definitely deserves a key position in the author list, but waive the authorship generously. We extend special thanks to Xiangyu Wu for sharing the most recent CLVisc code and providing helpful guidance for our simulations at various beam energies in later versions. We are also especially grateful to Longgang Pang for providing hydrodynamic profiles for our early numerical studies. We also thank Youyu Li, Xinxiang Liu, and Tan Luo for engaging in valuable discussions, digitizing experimental plots, and providing guidance on simulations. We thank Hengtong Ding, Xiaojian Du, Aiqiang Guo, Koichi Hattori, Min He, Che-Ming Ko, Yutie Liang, Shu Lin, Rapp Ralf, Kaijia Sun, Subhash Singha, Shuzhe Shi, Yifeng Sun, Aihong Tang, Biaogang Wu, Dilun Yang, Wenbin Zhao for valuable discussions. SL is supported by NSFC No. 12205090. Fundamental Research Funds for the central Universities. FL is supported by NSFC No. 12105129.